A survey about social media reported that 75​% of B2B marketers​ (marketers that focus primarily on attracting​ businesses) plan to increase their use of social​ media, as compared to 56​% of B2C marketers​ (marketers that primarily target​ consumers). The survey was based on 1,344 B2B marketers and 1,666 B2C marketers. The accompanying table summarizes the results. See Table

Increase Social Media Use   B2B Marketer   B2C Marketer   Total
Yes 1003 933 1936
No 341 725 1066
Total 1344 1658 3002

What is the probability that a randomly selected respondent plans to increase use of social​ media?​

What is the probability that a randomly selected respondent is a B2C​ marketer?​

What is the probability that a randomly selected respondent plans to increase use of social media or is a B2C​ marketer?​

Explain the difference in the results in​ (b) and​ (c).​

The two events in part​ (c) are collectively​ exhaustive, while the event in part​ (b) is not.

B.

The probability in part​ (b) is a joint​ probability, while the event in part​ (c) is a simple probability.

C.

The probability in part​ (b) is a marginal​ probability, while in part​ (c), the probability of an event or second event is calculated.

D.

The two events in part​ (c) are mutually​ exclusive, but part​ (b) has only one event.
Answer #1

a) P(increase use of social media) = 1936/3002=0.6449

b) P(P(B2C marketer) = 1658/3002=0.5523

c) P(randomly selected respondent plans to increase use of social media or is a B2C​ marketer) = (1003+933+725)/3002=0.8864

--------------------------

Explain the difference in the results in​ (b) and​ (c)

Answer:option c)

The probability in part​ (b) is a marginal​ probability, while in part​ (c), the probability of an event or second event is calculated